Calculate the boiling point elevation of an aqueous solution containing 36 grams of sucrose (C12H22O11) (MW=342g/mol) in 250 grams of water. The boiling point elevation constant for water is 0.512 degree Celsius per mole.


1
Expert's answer
2022-03-20T06:38:08-0400

ΔT = Kb m

m – molality

Moles of sucrose: 36 / 342 = 0.105

Mass of water: 250 g = 0.25 kg

So, molality = 0.105 / 0.25 = 0.42 mol/kg

ΔT = 0.512 x 0.42 = 0.2

Tboiling = 100 + 0.2 = 100.2 0C